OpenStudy (anonymous):

If w - 14 = 13, what is the value of w?

OpenStudy (anonymous):

Do you have an idea of the answer?

OpenStudy (anonymous):

w = 1 w = 18 w = 27 w = 32

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

i think that i need to add 14 on both sides

OpenStudy (anonymous):

Try it.

OpenStudy (anonymous):

w=27

OpenStudy (anonymous):

Correct.

OpenStudy (mynameisnemo):

w - 14 = 13 + 14 + 14 w=27 In order to figure out the value of w, you need to get w alone. In this case you need to cancel out the -14 by adding 14 to both sides (-14 + 14= 0 and 13+14= 27). w= 27

OpenStudy (anonymous):

Alot of people would choose w = 1 out of the simplicity. But all questions can't be easy.

OpenStudy (anonymous):

yea i got it

OpenStudy (anonymous):

w=27

OpenStudy (anonymous):

do you have any other questions?